How many angles are found in between 0 and 2π that satisfy the equation ItanxI = 5 ?


1
Expert's answer
2021-08-24T15:49:53-0400
"|\tan x|=5=>\tan x=\pm 5=>x=\pm\arctan 5+\pi n, n\in \Z"

"=>x=\pm\arctan 5+\pi n, n\in \Z"

For "[0, 2\pi]" we have


"x_1=\arctan 5, x_2=\pi-\arctan 5,"

"x_3=\arctan 5+\pi, x_4=2\pi-\arctan 5."

Four angles are found in between 0 and 2π that satisfy the equation ItanxI = 5.